# A Centralized Planning and Distributed Execution Method for Shape Filling with Homogeneous Mobile Robots

### v1

@@ 4 Proof of Correctness
    in part II, we establish the movement sequence $s[k]$, $t[k]$, $s_{r}[k]$ and $t_{r}[k]$ according to the orders established before. After that, we show that
      for an arbitrary epoch $T$ of the additive stage, the robot at $s[T]$ executing the edge-following movement stops at $t[T]$;
      for an arbitrary epoch $T$ of the subtractive stage, the robot at $s_{r}[T-N]$ executing the ribbon-following movement followed by the edge-following movement stops at $t_{r}[T-N]$.
    We conclude that the algorithm terminates with all lattice points in $S\setminus D$ occupied.
Due to the page limitation, only the outline for each part will be presented. For the complete proof, the reader could refer to appendix B for the proofs of part I and appendix C for the proofs of part II.
@@ 4.1 Part I
We start by placing the target shape $S$ in to the workspace. We assume that the placement is proper such that $S$ is placed entirely in the assembly half-plane $H_{1}$ and adjacent to the seed robots.
@@
Given a lattice point $x\in S$ with a positive hop count $k$, the neighborhood of $x$ can only be in one of the two types:
@@
Then we define undirected ribbons as the maximum continuous sequence of adjacent lattice points of the same hop count:
@@
  3. 3.
    every lattice point in $S$ has a positive hop count, that is,
    $(\forall x\in\{S\setminus D\}\cap L)h(x)>0$
  4. 4.
    $(\forall R\in\hat{S}$) $len(R)\geq 2$;
  5. 5.
    $(\exists!R\in\hat{S})g(R)=2\land len(R)=3$
@@
By satisfying the proper ribbon structure check in section 3, the ribbonization of $S$ is proper. We denote the ribbon with a gradient of $2$ as $R_{0}$. This ribbon serves as the root of the tree of ribbons:
@@
    $(\forall R_{i},R_{j}\in\hat{S})R_{i}<R_{j}\iff R_{i}\rightarrow R_{j}$
@@ 4.2 Part II
We prove the correctness of the algorithm in two steps: for the additive stage, we show that the robot at $s[T]$ executing the edge-following movement stops at $t[T]$; for the subtractive stage, we show that the robot at $s_{r}[T-N]$, executing the ribbon-following movement followed by the edge-following movement, stops at $t_{r}[T-N]$.
@@ 4.2.1 Additive Stage
By induction on the perimeter of the swarm in each epoch, together with the boundary smoothness condition (required by the proper ribbonization), we will show that there is a unique and unobstructed path for robot $u_{T}$ when it executes the edge-following movement and $t[T]$ is on this unique path.
@@
Since each active robot $u_{T}$ executing the edge-following algorithm has an unobstructed path from $s[T]$ and $t[T]$, and $u_{T}$ is localizable at $s[T]$ and $t[T]$, we conclude that:
Theorem 5.
In each epoch $T$ with $1\leq T\leq N$, the idle robot located at $s[T]$ starts to move and stops at $t[T]$.
@@ 4.2.2 Subtractive Stage
To facilitate the proof, we define the merging point $\mathbb{m}$ for each ribbon as the first lattice point a ribbon-following robot encounters after exiting the shape $S$. Next, we investigate the perimeter of the swarm after $u_{N}$ stops, and show that the merging point is in front of all the other points on the perimeter contributed by the same ribbon. This property ensures the robot exiting $S$ follows the perimeter of the swarm by executing the edge-following algorithm. The path from $s_{r}[T-N]$ to $t_{r}[T-N]$ is divided by the merging point into the interior path and the exterior path:
@@
In each epoch $T$ with $N+1\leq T\leq 2N$, robot $s_{T}$ is localizable at lattice point $s_{r}[T-N]$, $t_{r}[T-N]$ and the merging point $m$ (if the robot exits the shape $S$).
Since each active robot $u_{T}$ executing the ribbon-following and the edge-following algorithm has an unobstructed path from $s_{r}[T-N]$ to $t_{r}[T-N]$, and $u_{T}$ is localizable at $s_{r}[T-N]$, $t_{r}[T-N]$ and the merging point, we conclude that:
@@
In each epoch $T$ with $N+1\leq T\leq 2N$, the robot $s_{T}$ locating at $s_{r}[T-N]$ stops at $t_{r}[T-N]$.
Finally, by theorem 5 and theorem 8, together with proposition 4 and proposition 7, we conclude that:
Theorem 9.
@@
    For each prefix subsequence $(a_{1},b_{1},1),\ldots,(a_{k},b_{k},k)$, $1\leq k\leq 2N$, let $A_{k}=\{a_{1},\ldots,a_{k-1}\}$, $B_{k}=\{b_{1},\ldots,b_{k-1}\}$, and $\overline{B_{k}}=B_{k}$ since the add-subtract algorithm enforce at most one moving robot at any time. Then the starting position of the $k$th robot $a_{k}\in I\oplus\overline{B_{k}}\circleddash A_{k}$ and the robot can only relay on robots centered at $I\oplus\overline{B_{k}}\circleddash A_{k}\circleddash\{a_{k}\}$ to locate its position.
@@ 5.1 Limitation of the Algorithm
In fact, the add-subtract algorithm works for all connected shape $S$ which can be discretized into a proper ribbon structure. Roughly speaking, a "proper" ribbon structure of $S$ must ensure that each ribbon has a minimum length of two, and the lattice points on the boundary of $S$ must exhibit certain smoothness. Fig. 15 show the simulation of the failure of forming a hollow $C$-shape, which lacks a smooth boundary. Notice that the robot $u_{142}$ (indicated by the red arrow) gets stuck at the inner boundary.
@@ 5.2 Real World Application
Suppose we aim to use UAVs to monitor the flow of people in the Punggol park, a community park located in the Hougang area of northeastern Singapore. Noticing that the park encompasses a lake, and no UAVs are required to be deployed over the lake. In this scenario, the add-subtract algorithm provides an efficient way to deploy the UAVs:
@@
6 Robustness Analysis
To handle failure of units, additional mechanism is implemented. First, we shall extend the activation sequence $s$ to contain all the lattice points that has been occupied by the idle robots at the initial configuration. Suppose the malfunctioning robot is not able to move but is still able to communicate with neighboring robots. For epoch $T$, with the malfunctioning robot $u_{T}$, the following cases are considered:
  1. 1.
    $T\leq N$ and the malfunctioning robot stops outside the desired shape, then the robot broadcast a message to trigger all the other robots to shift the unexecuted activation subsequence forward by $1$ element, i.e., $s_{[T:N]}=s_{[T+1:N+1]}$.
  2. 2.
    $T\leq N$ and the malfunctioning robot stops inside the desired shape at lattice position $x$, then the robot broadcast a message to trigger all the other robots to replace the element in $t$ of value $x$ with $t[T]$.
  3. 3.
    $T>N$ and the malfunctioning robot stops outside the desired shape at lattice position $x$, then the robot broadcast a message to trigger all the other robots to shift the unexecuted reassembly subsequence backward by $1$ element, i.e., ${t_{r}}_{[T-N+1:2N]}={t_{r}}_{[T-N:2N-1]}$ and start the next epoch.
Notice that in either case, the movement sequences are altered in such the way that preserves the tree order and idle tree order. As a result, the above mechanism can effectively handle the failure of any robot, provided that the malfunctioning robot does not block the way of subsequent moving robots. In other word, with the additional failure-handling mechanism, the entire algorithm only fails when:
  1. 1.
    the malfunctioning robot is in the subtractive stage ($T>N$) and the malfunctioning robot stops in $S$; or
  2. 2.
    the malfunctioning robot stops outside $S$ at some lattice point $x$, and $x$ decreases the boundary smoothness such that $S\cup\{x\}$ does not satisfy the boundary smoothness requirement in section LABEL:sect-probDescription.
@@ 7 Conclusion and Future Works
8 Acknowledgment
ChatGPT [15] has been used to improve the syntax and grammar for the abstract and introduction of the manuscript.
@@ Appendix A Algorithms
$s[k],t[k],s_{r}[k],t_{r}[k]$, rotation radius $d$, synchronized time $t$ [^1] Time $t$ is synchronized between all robots of the swarm at the start of the algorithm, and updated locally while executing the algorithm. Thus, for simplicity, we assume that all the robots of the swarms shares a synchronized clock. and warm-up time $t_{warmup}$
@@ Appendix B Ribbon and its Properties
In appendix B, we discuss the ribbonization procedure: ribbonization essentially partitions the shape $S$ (represented by a set of discrete lattice points) as layers stacking on one another. Each layer is captured by a set of lattice points in $S$, and each set of points forms a line graph if we impose an edge between adjacent lattice points. We call the path graphs induced by such sets (or equivalently, the order list of vertex points of the path graphs) "ribbons". There also exists a tree structure between ribbons and this property is essential when we define the movement sequence in section C.1 and section C.5.
For generalization purposes, we start with an arbitrary shape $P_{0}$ that is homeomorphic to an open disk. Suppose its corresponding discrete representation is $P$. To partition the lattice points in $P$, we first introduce the shape $P$ into the workplace with the underlying hexagonal lattice, certain requirements are imposed to ensure a "proper" placement. Then, we introduce the hop count defined on every lattice point in $P$ that denotes the distance from the lattice point to the origin. This hop count value depends on the shape of $P$ in the sense that it represents the length of the shortest path within $P$. The properties of hop count are discussed in section B.1. We introduce ribbons as adjacent lattice points in $P$ with the same hop count value. Ribbons are defined in section B.2 and the ribbonization of shape $P$ is investigated in section B.3. Lastly, we consider the two special cases for $P$: in section B.4, we let $P$ be the user-defined shape $S$ and we obtain ribbons of $S$; in section B.5, we let $P$ to be the set of lattice points in $H_{2}$ and we obtain the idle ribbons.
@@ B.1 Hop count
For any lattice point $x$ with a positive hop count $k$, $v$ cannot have two none-adjacent neighbors with hop count $k-1$.
@@
If $v$ is surrounded by $path_{u}$ and $path_{w}$, the straight line path starts from $u$ in the direction of $v$ and the straight line path starts from $w$ in the direction of $v$ must intersect $path_{u}$ or $path_{w}$. It contradicts $path_{u}$ being the shortest path or $path_{w}$ being the shortest path, respectively.
@@
    Type $A$: $x$ is adjacent to one lattice point $y$ with hop count $k-1$. The lattice point(s) in $P$ that are adjacent to both $x$ and $y$ has the hop count equal to $k$. The remaining neighbors in $P$, if any, have hop count equal to $k+1$.
@@
  • case 1: $x$ has only one neighbor $y$ with hop count $k-1$. We shall inspect the six neighbors of $x$. Denote the two lattice points adjacent to both $x$ and $y$ as $a$ and $b$. If $a,b$ inside $P$, then they could have hop count equal to $k$ or $k+1$ by proposition 1. Since they are adjacent to $y$ with $h(y)=k-1$, $a,b$ must have hop count $k$. Denote the lattice point adjacent to $a,x$ as $c$ and the lattice point adjacent to $b,x$ as $d$. Suppose $c,d$ locate in $P$, then by proposition 5, $h(c)\neq k$ and $h(d)\neq k$ so $h(c)=h(d)=k+1$. Denote the last lattice point as $e$ and assume $e\in P$. By applying proposition 3 on lattice point $c$, we deduce that $h(e)\neq k$, so $h(e)=k+1$. We conclude that the neighborhood of $x$ in case 1 is precisely the type A neighborhood of $x$.
@@ B.3 Ribbons Induced by $P$
    $R_{i}$ is a descendant ribbon of $R_{j}$ if there is a chain $(R_{i},R_{i+1},\ldots,R_{j-1},R_{j})$ such that $R_{i}$ is a child ribbon of $R_{i-1}$, …, $R_{j-1}$ is a child ribbon of $R_{j}$. We also denote $R_{j}$ as an ancestor ribbon of $R_{i}$.
@@ B.4 Ribbonization of $S$
Definition 14.
A proper ribbonization of $S$ satisfies the following properties:
  1. 1.
    all conditions in the proper ribbonization of $S$ as in definition 9; and
  2. 2.
    every lattice point on the boundary of $S$ is not in $D$, that is
@@
Note that:
  1. 1.
    definition 14.2 ensures the hole $D$ could be correctly captured. In other words, no lattice point in the hole $D$ is on the boundary of $S$.
@@ B.5 Ribbonization of Half-plane $H_{2}$
Observe that any lattice point in $H_{2}$ belongs to some idle ribbon and the idle ribbons preserve the properties discussed previously.
@@ B.6 Ordering Induced by the Ribbon Structure
The idle order is a strict total order of the lattice points in $H_{2}$ such that for all $x,y\in H_{2}$, $x<y$ iff
@@
    $h(x)=h(y)$ and $x$ is behind $y$ with respect to the idle ribbon.
@@
    Given a ribbon $R$, the ribbon order is a strict total order of the lattice points of $R$ such that for all $x,y\in V(R)$, $x<y$ iff $x$ is in front of $y$.
@@
    Given a ribbonization tree $T(\hat{S})$, the tree order of a ribbonization tree $T$ is the strict partial order on the set of ribbons $\hat{S}$ such that $\forall R_{i},R_{j}\in\hat{S}$, $R_{i}<R_{j}$ iff the directed path from the root $R_{0}$ to $R_{j}$ passes through $R_{i}$. The complete tree order is a strict total order that extends the tree order [^14] A partial order extends to a total order by the Szpilrajn extension theorem..

### v3

@@ 4 Proof of Correctness
    in part II, we establish the movement sequences $s[k]$, $t[k]$, $s_{r}[k]$ and $t_{r}[k]$ according to the orders established before. After that, we show that:
      for an arbitrary epoch $T$ of the additive stage, the path from $s[T]$ to $t[T]$ is unobstructed, and the moving robot can localize itself whenever necessary;
      for an arbitrary epoch $T$ of the subtractive stage, the path from $s_{r}[T-N]$ to $t_{r}[T-N]$ is unobstructed, and the moving robot can localize itself whenever necessary.
    Notice that the activation sequence $s$, the reactivation sequence $s_{r}$, the assembly sequence $t$ and the re-assembly sequence $t_{r}$ uniquely defined the sequence of tuples $(s_{1},t_{1},1),\ldots,(s_{N},t_{N},N),({s_{r}}_{1},{t_{r}}_{1},N+1),\ldots,({s_{r}}_{N},{t_{r}}_{N},2N)$. We conclude that the sequence of tuples satisfies the problem statement.
Due to the page limit, only the outline for each part will be presented. For the complete proof, the reader could refer to appendix B for the proofs of part I and appendix C for the proofs of part II.
@@ 4.1 Part I
We start by placing the target shape $S$ in the workspace. We assume that the placement is proper such that $S$ is placed entirely in the assembly half-plane $H_{1}$ and adjacent to the seed robots.
@@
Given a lattice point $x\in S$ with a positive hop count $k$, the neighborhood of $x$ can only be one of the two types:
@@
Then, we define undirected ribbons as the maximum continuous sequence of adjacent lattice points of the same hop count:
@@
    every lattice point in $S$ has a positive hop count, that is,
    $(\forall x\in\{S\setminus D\}\cap L)h(x)>0$
  3. 3.
    $(\forall R\in\hat{S}$) $len(R)\geq 2$;
  4. 4.
    $(\exists!R\in\hat{S})g(R)=2\land len(R)=3$
  5. 5.
@@
Notice that, by satisfying the proper lattice representation requirements in section 2.1 and the proper ribbon structure requirements in section 3.1, the ribbonization of $S$ is proper. We denote the ribbon with a gradient of $2$ as $R_{0}$. This ribbon serves as the root of the tree of ribbons.
@@
    $(\forall R_{i},R_{j}\in\hat{S})R_{i}<R_{j}\iff R_{i}\twoheadrightarrow R_{j}$
    The complete tree order is a strict total order that extends the tree order [^10] A partial order extends to a total order by the Szpilrajn extension theorem..
@@ 4.2 Part II
We prove the correctness of the algorithm in two steps: for the additive stage, we show that the path between $s[T]$ and $t[T]$ is unobstructed, and the active robot could localize itself using non-active robots at $s[T]$ and $t[T]$; for the subtractive stage, we show that the path between $s_{r}[T-N]$ and $t_{r}[T-N]$ is unobstructed, and the active robot could localize itself using non-active robots at $s_{r}[T-N]$, $t_{r}[T-N]$ and the merging point (defined shortly).
@@ 4.2.1 Additive Stage
By induction on the perimeter of the MRS in each additive epoch, together with the boundary smoothness condition (required by the proper ribbonization), we show that there is a unique and unobstructed path for robot $u_{T}$ when it executes the edge-following movement and $t[T]$ is on this unique path.
@@ 4.2.2 Subtractive Stage
To facilitate the proof, we define the merging point $\mathbf{m}$ for each ribbon as the first lattice point a ribbon-following robot encounters after exiting the shape $S$. Next, we investigate the perimeter of the swarm after $u_{N}$ stops, and show that the merging point is in front of all the other points on the perimeter contributed by the same ribbon. This property ensures that the robot exiting $S$ follows the perimeter of the swarm by executing the edge-following algorithm. The path from $s_{r}[T-N]$ to $t_{r}[T-N]$ is divided by the merging point into the interior path and the exterior path:
@@
In each epoch $T$ with $N+1\leq T\leq 2N$, robot $s_{T}$ is localizable at lattice point $s_{r}[T-N]$, $t_{r}[T-N]$ and the merging point $m$.
Finally, by proposition 3, proposition 4, proposition 5 and proposition 6, we conclude that:
@@
Given sufficient robots (with an initial set of robots $B$ whose position coordinates are known in advance) and the user-defined shape $S\setminus D$. The add-subtract algorithm with the movement sequence of tuples $(s_{1},t_{1},1),\ldots,(s_{N},t_{N},N),({s_{r}}_{1},{t_{r}}_{1},N+1),\ldots,({s_{r}}_{N},{t_{r}}_{N},2N)$ satisfies:
@@
    For each prefix subsequence $(a_{1},b_{1},1),\ldots,(a_{k},b_{k},k)$, $1\leq k\leq 2N$, let $A_{k}=\{a_{1},\ldots,a_{k-1}\}$, $B_{k}=\{b_{1},\ldots,b_{k-1}\}$, and $\overline{B_{k}}=B_{k}$. Then the starting position of the $k$th robot $a_{k}\in I\oplus\overline{B_{k}}\circleddash A_{k}$ and the robot can only relay on robots centered at $I\oplus\overline{B_{k}}\circleddash A_{k}\circleddash\{a_{k}\}$ to locate its position before stopping at $b_{k}$.
@@ 5.1 Limitation of the Algorithm
In fact, the add-subtract algorithm works for all connected shapes $S$ which can be discretized into a proper ribbon structure. Roughly speaking, a "proper" ribbon structure of $S$ must ensure that each ribbon has a minimum length of two, and the lattice points on the boundary of $S$ must exhibit certain smoothness. Fig. 15 shows the simulation of the failure of forming a hollow $C$-shape, which lacks a smooth boundary. Notice that the robot $u_{142}$ (indicated by the red arrow) gets stuck at the inner boundary.
@@ 5.2 Real World Application
Suppose we aim to use UAVs to monitor the flow of people in Punggol Park, a community park located in the Hougang area of northeastern Singapore. Noticing that the park encompasses a lake, and no UAVs are required to be deployed over the lake. In this scenario, the add-subtract algorithm provides an efficient way to deploy the UAVs:
@@ Appendix A Algorithms
$s[k],t[k],s_{r}[k],t_{r}[k]$, rotation radius $d$, synchronized time $t$ [^1] Time $t$ is synchronized between all robots of the swarm at the start of the algorithm and updated locally while executing the algorithm. Thus, for simplicity, we assume that all the robots of the swarms share a synchronized clock. and warm-up time $t_{warmup}$
@@ Appendix B Ribbon and its Properties
In appendix B, we discuss the ribbonization procedure: ribbonization essentially partitions the shape $S$ (represented by a set of discrete lattice points) as layers stacking on one another. Each layer is captured by a set of lattice points in $S$, and each set of points forms a line graph if we impose an edge between adjacent lattice points. We call the path graphs induced by such sets (or equivalently, the order list of vertex points of the path graphs) "ribbons". There also exists a tree structure between ribbons, and this property is essential when we define the movement sequence in section C.1 and section C.5.
For generalization purposes, we start with an arbitrary shape $P_{0}$ that is homeomorphic to an open disk. Suppose its corresponding discrete representation is $P$. To partition the lattice points in $P$, we first introduce the shape $P$ into the workplace with the underlying hexagonal lattice, and certain requirements are imposed to ensure a "proper" placement. Then, we introduce the hop count defined on every lattice point in $P$ that denotes the distance from the lattice point to the origin. This hop count value depends on the shape of $P$ in the sense that it represents the length of the shortest path within $P$. The properties of hop count are discussed in section B.1. We introduce ribbons as adjacent lattice points in $P$ with the same hop count value. Ribbons are defined in section B.2 and the ribbonization of shape $P$ is investigated in section B.3. Lastly, we consider the two special cases for $P$: in section B.4, we let $P$ be the user-defined shape $S$ and we obtain ribbons of $S$; in section B.5, we let $P$ to be the set of lattice points in $H_{2}$ and we obtain the idle ribbons.
@@ B.1 Hop count
For any lattice point $x$ with a positive hop count $k$, $v$ cannot have two non-adjacent neighbors with hop count $k-1$.
@@
If $v$ is surrounded by $path_{u}$ and $path_{w}$, the straight line path starts from $u$ in the direction of $v$, and the straight line path starts from $w$ in the direction of $v$ must intersect $path_{u}$ or $path_{w}$. It contradicts $path_{u}$ being the shortest path or $path_{w}$ being the shortest path, respectively.
@@
    Type $A$: $x$ is adjacent to one lattice point $y$ with hop count $k-1$. The lattice point(s) in $P$ that are adjacent to both $x$ and $y$ has the hop count equal to $k$. The remaining neighbors in $P$, if any, have the hop count equal to $k+1$.
@@
  • case 1: $x$ has only one neighbor $y$ with hop count $k-1$. We shall inspect the six neighbors of $x$. Denote the two lattice points adjacent to both $x$ and $y$ as $a$ and $b$. If $a,b$ are inside $P$, then they could have the hop count equal to $k$ or $k+1$ by proposition 1. Since they are adjacent to $y$ with $h(y)=k-1$, $a,b$ must have hop count $k$. Denote the lattice point adjacent to $a,x$ as $c$ and the lattice point adjacent to $b,x$ as $d$. Suppose $c,d$ locate in $P$, then by proposition 5, $h(c)\neq k$ and $h(d)\neq k$ so $h(c)=h(d)=k+1$. Denote the last lattice point as $e$ and assume $e\in P$. By applying proposition 3 on lattice point $c$, we deduce that $h(e)\neq k$, so $h(e)=k+1$. We conclude that the neighborhood of $x$ in case 1 is precisely the type A neighborhood of $x$.
@@ B.3 Ribbons Induced by $P$
    $R_{i}$ is a descendant ribbon of $R_{j}$ if there is a chain $(R_{i},R_{i+1},\ldots,R_{j-1},R_{j})$ such that $R_{i}$ is a child ribbon of $R_{i-1}$, …, $R_{j-1}$ is a child ribbon of $R_{j}$. We also denote $R_{j}$ as an ancestor ribbon of $R_{i}$, denoted as $R_{j}\twoheadrightarrow R_{i}$.
@@ B.4 Ribbonization of $S$
Recall that every lattice point on the boundary of $S$ is not in $D$, that is
@@ B.5 Ribbonization of Half-plane $H_{2}$
Observe that any lattice point in $H_{2}$ belongs to some idle ribbon, and the idle ribbons follow the properties discussed previously.
@@ B.6 Ordering Induced by the Ribbon Structure
    $(\forall IR_{i},IR_{j}\in\hat{H_{2}})IR_{i}<IR_{j}\iff IR_{i}\twoheadrightarrow IR_{j}$
@@
    $(\forall x,y\in V(R))x<y\iff y\rightarrow x$
@@
    $(\forall R_{i},R_{j}\in\hat{S})R_{i}<R_{j}\iff R_{i}\twoheadrightarrow R_{j}$
    The complete tree order is a strict total order that extends the tree order [^15] A partial order extends to a total order by the Szpilrajn extension theorem..